OpenStudy (anonymous):

The equation x^2+y^2=25 defines a circle with center at the origin and radius 5. The line y=x-1 passes through the circle. Using the substitution method, find the point(s) at which the circle and the line intersect. a. (4,-3) and (-3,4) b. (3,4) and (-3,-4) c. (4,3) and (-3,-4) d. No solution

OpenStudy (anonymous):

see he gave a line " y = x -1" and he also said it passes through circle, so just substitute that y value ( which is x -1 right ? ,from line eq), in circle equation and solve that quadratic eqaution for x and then finally subtitute those two x values you got by solving that quadratic eqn in the line equation y = x -1 to get the corresponding y values

OpenStudy (anonymous):

sorry had to type that much theory , but i hope you get what i'm trying to say :)

OpenStudy (anonymous):

Ok, I think I got it. I got (4,-3) and (-3,4) which is a. Is that right?

OpenStudy (anonymous):

what are the values of x you got?

OpenStudy (anonymous):

because those points you got should also satisfy that line y = x -1 but your points are not satisfying the line?

OpenStudy (anonymous):

I got x=4 and x=-3